OpenStudy (anonymous):

h=-3cos(5π/3)t+3.5 Where t is the time measured in seconds. a. What is the highest point reached? b. What is lowest point reached? c. What is the period of the model? d. According to the model, find the height after 25 seconds.

OpenStudy (anonymous):

biggest cosine can be is 1 so biggest -3 cosine can by is 3 and so biggest whole thing can be is 6.5 smallest cosine can be is -1, so smallest -3 cosine can be is -3 so smallest whole thing can be is .5

OpenStudy (anonymous):

period is \[\frac{2\pi}{b}\] in this case \[\frac{2\pi}{\frac{5\pi}{3}}=\frac{6}{5}\]

OpenStudy (anonymous):

for the last one replace t by 25